Suppose that a child's grade point average is directly proportional to that child's hours of outdoor play. Last semester, nine-year-old Jamie played outdoors an average of 20 hours per week, and a great point average was 3.25. So far this semester, she has only managed to play outdoors 12 hours/week. What grade point average for this produce?
1
Expert's answer
2017-09-25T12:42:06-0400
If a child's grade point average is directly proportional to that child's hours of outdoor play, then grade point average for this produce is:
Grade/12 = 3.25/20,
Grade = 12*3.25/20 = 1.95.